# The NOvA Collaboration

### v1

@@ IV COSMIC MUON DATA IN THE NOvA NEAR DETECTOR
Figure 6: Logarithm (base 10) of time difference between $i$ and $(i+1)$th multiple-muon events from one day of data. The distribution is fitted with an Erlang distribution in the range 1.5 to 4.7. Error bars represent statistical uncertainties.
Figure 7: $\chi^{2}$/NDF distribution of Erlang fit for multiple-muon data in weekly binning.
V Temperature data
Atmospheric temperature data are taken from the European Centre for Medium-Range Weather Forecasts (ECMWF) Dee et al. (2011). For this analysis, we utilize temperature data from ERA5 Hersbach et al. (2020), the fifth-generation atmospheric reanalysis produced by ECMWF. They provide temperature readings at 37 pressure levels, distributed nonuniformly from 1 to 1000 hPa. For comparison with the NOvA ND data, temperatures are extracted from ECMWF within the grid area ranging from (41N, 87W) to (42.5N, 88.5W).
$T_{e\!f\!f}=\frac{\sum_{i=1}^{n}{T(P_{i})[W_{\pi}(P_{i})+W_{K}(P_{i})]\Delta P_{i}}}{\sum_{i=1}^{n}[W_{\pi}(P_{i})+W_{K}(P_{i})]\Delta P_{i}}.$ (2)
Here, $T(P_{i})$ is the temperature at pressure level $i$ ($P_{i}$), and $W_{K}(P_{i})$ and $W_{\pi}(P_{i})$ are the weights responsible for the contribution as a function of pressure (altitude) for those $K$ and $\pi$ decays that make muons capable of reaching the NOvA ND. The effective temperature is calculated using cross sections and lifetime values from Grashorn et al. (2010), along with temperature values obtained at 37 pressure levels from ECMWF. Figure 8 shows the variation of ECMWF temperature data of the atmosphere over the Fermilab area, averaged over the four years 2018–2022, as a function of altitude and pressure. It also shows the total weight $W_{\pi}+W_{K}$ used in Equation 2.
Figure 8: Variation of average temperature with altitude and total weight function used to calculate effective temperature.
VI Cosmic muon simulation
Cosmic muons are simulated with CORSIKA Heck et al. v7.7410 using GHEISHA as a low-energy hadronic interaction model and QGSJET as a high-energy hadronic interaction model. The energy of primary particles is chosen to be between $200$ and $10^{6}$ GeV following a power law with exponent 2.7-2.7 Navas et al. (2024). Each muon in a multiple-muon event is required to have an energy above $50\text{\,}\mathrm{G}\mathrm{e}\mathrm{V}$, which is necessary to reach the NOvA ND vertically. A muon coming at a high zenith angle has to pass through a larger overburden and hence requires more energy to reach underground detector compared to vertically incident muons. This is taken care of by calculating the energy loss using Equation 3. Cosmic showers are simulated using the Earth’s magnetic field for the Fermilab location (41.84N, 88.25W).
The energy loss of muons due to the overburden above the ND is calculated using the relation
$E_{\mu}=(E_{\mu,0}+\epsilon)e^{-bd\rho}-\epsilon,$ (3)
where $E_{\mu,0}$ and $E_{\mu}$ denote the initial and final energies of the muons, respectively, as they traverse a distance $d$ through the underground rock with density $\rho$. The parameters $\epsilon$ and $b$ are material-dependent constants, with values of $500\text{\,}\mathrm{G}\mathrm{e}\mathrm{V}$ and $4\times 10^{-6}~\text{cm}^{2}/\text{g}$, respectively, for standard rock Zyla et al. (2020). Muons that retain nonzero energy after passing through the overburden are considered for the rate calculation below.
CORSIKA models the change in atmospheric density with altitude using a five-layer structure that extends from the ground to the top of the atmosphere ($100\text{\,}\mathrm{k}\mathrm{m}$). The density in the bottom four layers decreases exponentially with altitude, described in terms of the atmospheric mass overburden $X(h)$ and height $h$ by
$X(h)=a_{i}+b_{i}e^{-\frac{h}{c_{i}}},$ (4)
where $i$ ranges from 1 to 4. As the atmosphere becomes exceedingly thin in the top layer, the mass overburden decreases linearly with height for the 5th layer:
$X(h)=a_{5}-b_{5}h/c_{5}.$ (5)
CORSIKA provides default atmospheric profiles for some selected locations but does not support the Fermilab site. To address this, we extracted vertical atmospheric profile data during 2017 from the Global Data Assimilation System (GDAS) National Centers for Environmental Prediction (2024) (NCEP) and matched CORSIKA’s exponential layer model. The site-specific atmospheric coefficients ($a_{i}$, $b_{i}$, and $c_{i}$) are used as input parameters to the simulation. Proton showers are generated using simulation parameters described earlier in this section and atmospheric parameters extracted from GDAS. We found good agreement between the simulation and the data collected by the NOvA experiment in the ND, as shown in  Figure 5.
In the study of the seasonal effect, the single and multiple-muon event rates are calculated by using the events on a daily basis, and later binned by month. The passage of events through NOvA ND is simulated using GEANT Agostinelli et al. (2003); Allison et al. (2006, 2016).
After the generation of showers, cosmic muons from each generated shower are spread uniformly (randomized) over an area from $32\text{\,}\mathrm{m}$ to $32\text{\,}\mathrm{m}$ along the $x$-axis (horizontal) and $30\text{\,}\mathrm{m}$ to $46\text{\,}\mathrm{m}$ along the $z$-axis (along the neutrino beam) of the NOvA coordinate system. For multiple-muon events, the first muon is randomized uniformly over the $xz$ surface, and the remaining muons are placed on the $xz$ surface without changing the distance between muons and orientation given by CORSIKA. The ND’s surface area spans from $2\text{\,}\mathrm{m}$ to $2\text{\,}\mathrm{m}$ along the $x$-axis and from 0 to $16\text{\,}\mathrm{m}$ along the $z$-axis; the extended area for randomization is used to allow muons to hit the side of the detector. This choice allows us to include the core of 99.3% of all cosmic muons that hit the detector, and is much larger than the actual detector. Showers that produce at least one muon in the ND proceed to the detector simulation. Neutrino-induced upward-going muons are not considered for this analysis.
We studied the seasonal results for muons from two distinct stages of the simulation. We show results for all muons in a CORSIKA event that have sufficient energy to reach the underground detector. We refer to these results as predictions for an infinite detector, for which there was no detector simulation. We also used the CORSIKA output to simulate events in the NOvA ND as described in the previous paragraph. In this case, the seasonal results were obtained with the same software as used for NOvA ND data Acero et al. (2020). It was the marked difference in the results before and after the detector simulation that supported the hypothesis that a geometry effect was involved.
We do not claim that our implementation of CORSIKA accurately represents the altitude and temperature distributions in the summer and winter. In fact, the magnitude of the seasonal modulation in our CORSIKA simulation does not match the expected magnitude based on the well-established theoretical and experimental understanding of single-muon seasonal behavior Grashorn et al. (2010). However, we do expect the density profile in the atmosphere to provide a fairly good basis for the summer/winter differences regarding the hadronic showers and the competition between hadron interaction and decay in the upper parts of the atmosphere. Our use of the simulation in this analysis relies only on a comparison of the seasonal phases and magnitude for single- and multiple-muon events in a detector of fixed size.
VII Fits to Seasonal Variations
The simulated multiple-muon rate for an infinite detector shows a summer maximum as displayed in Figure 9, opposite that seen in multiple-muon measurements. When the simulated infinite detector events were passed through the finite ND, the time of the seasonal maximum moved from summer to winter, as shown in Figure 10. This result from our simulation is consistent with previous data observations (NOvA Acero et al. (2019) and MINOS Adamson et al. (2015)).
Figure 9: Seasonal variation of multiple-muon rate for an infinite detector at the depth of the NOvA ND from CORSIKA initiated by primary proton cosmic rays. The solid curve shows a cosine fit.
Figure 10: Seasonal variation of multiple-muon rate at the NOvA ND (finite detector) from CORSIKA initiated by primary proton cosmic rays. The solid curve represents a cosine fit of the variation. Error bars represent statistical uncertainties.
@@ VII Fits to Seasonal Variations
Figure 11 shows the multiple-muon event rate seen in the four-year (May 2018 to April 2022) NOvA analysis sample discussed above, as well as the effective temperature. Both the data and the temperature are presented in weekly bins. The percentage variation is calculated using
@@
The muon rate in the multiple-muon data displays a seasonal trend that is out of phase with the effective atmospheric temperature. Although periodic, there is no a priori reason to expect a cosine shape. However, it does serve to extract the phase of the magnitude maxima, so the percentage variations of multiple-muon rate and effective temperature are fitted with cosine curves:
@@
where $V_{0}$ is the average rate, $V$ is the magnitude, $t_{year}$ is the time period (fixed at one year) and $\phi$ is the phase. The fitted parameters are shown in Table 1. The multiple-muon data vary by $\pm 2\%$, and have the opposite phase as the temperature.
Table 1: Best fitted parameters for the cosine fit of percentage variation of the multiple-muon data and effective temperature
@@
VIII Other Seasonal Effects
In addition to the aforementioned altitude–geometry effect, the MINOS paper Adamson et al. (2015) discussed three more ideas to explain the observed seasonal pattern of multiple-muon event rate: (1) hadronic dimuon decays, (2) a temperature effect in layers of the atmosphere, and (3) an anticorrelation of primary and secondary meson decays. Another possible explanation is muon decay. We next discuss each in turn.
(1) Although the decay probability of pions increases in the summer, the decay probabilities of other hadrons, such as $\eta$ and $\rho$ mesons that decay into dimuons, change negligibly. However, due to the small dimuon branching ratio of these mesons (branching ratios for $\rho\rightarrow\mu^{+}\mu^{-}$ and $\eta\rightarrow\mu^{+}\mu^{-}\gamma$ are $4.6\times 10^{-5}$ and $3.1\times 10^{-4}$, respectively) hadronic dimuon decays are not sufficient to explain the opposite behavior in multiple-muon event rate compared to single muons.
(2) The temperature effect mentioned in Adamson et al. (2015) implies that there could be a seasonal variation dependent on altitude, differing for single and multiple-muon event rates. In the MINOS paper Adamson et al. (2015), the temperature data were examined to see any significant temperature changes, leading to changes in rate, at a particular altitude, but no such effect was observed.
(3) In the MINOS paper Adamson et al. (2015), the most plausible explanation for the observed phenomenon was considered to be the anticorrelation between the primary and secondary mesons. According to this hypothesis, the presence of multiple muons may arise not solely from the decay of mesons generated in the initial interaction, but also from mesons produced in subsequent interactions involving secondary mesons and air molecules. The higher probability of meson interactions in winter was posited as a contributing factor, leading to an increased number of mesons that, upon decay, yield muons and consequently contribute to the occurrence of more multiple-muon events. If this were true, the multiple-muon rate should exibit a winter maximum in the infinite detector. This explanation is contradicted by our simulation results; in Figure 9 we observed a summer maximum.
(4) Another potential explanation for the observed seasonal pattern of multiple muons is muon decay. Due to relativistic effects, this phenomenon is most pronounced at low muon energies, as observed by detectors at or near the surface. The seasonal effect of muon decay for various muon thresholds can be found in Table 2. Surface detectors such as the NOvA far detector Acero et al. (2021), DECOR Yurina et al. (2019); Tolkacheva et al. (2011), and GRAPES Arunbabu et al. (2017), all with muon thresholds near 5 GeV, can exhibit a 0.6% excess of events during the winter. In contrast, detectors such as the NOvA ND and MINOS ND, which have muon thresholds of 50 GeV, would only show a smaller 0.09% excess in winter. This effect is even smaller for the MINOS FD and MACRO.
@@ VIII Other Seasonal Effects
Table 2: A prediction of the percentage of muons that reach sea level before they decay as a function of muon energy. $E_{\mu}$ is the muon energy, $N_{0}$ is the initial number of muons, and $N$ is the number of muons that survive. The summer/winter altitude difference is assumed to be 2%.
@@
IX Conclusion
@@
The altitude–geometry effect is the most relevant explanation for the opposite seasonal behaviors between single and multiple-muon events. Evidence supporting this conclusion comes from a comparison between the typical muon separation distribution as a function of underground detector depth vis-a-vis the detector size, as well as a CORSIKA simulation, which demonstrates a summer maximum for multiple-muon events in an infinite detector. However, for a finite-sized detector, such as the NOvA ND, the simulation shows a winter maximum for multiple-muon events. Other proposed solutions would have resulted in a winter maximum in the CORSIKA simulation for an infinite detector. Additionally, the explanation of muon decay was found to be too small to account for observations of $50\text{\,}\mathrm{G}\mathrm{e}\mathrm{V}$ muons. The altitude–geometry effect was found to be consistent with observations from the MINOS FD, MINOS ND, NOvA ND, MACRO, and NOvA FD. All other considered explanations were excluded on multiple grounds.

### v2

@@ V COSMIC MUON DATA IN THE NOvA NEAR DETECTOR
Figure 6: Logarithm (base 10) of time difference (in ms) between $i$ and $(i+1)$th multiple-muon events from one day of data. The distribution is fitted with an Erlang distribution in the range 1.5 to 4.7, yielding an observed multiple muon rate of $0.144\pm 0.003$ Hz for that day. Selected noise events that do not match the hypothesis of a cosmic ray multiple muon of that rate are seen on the left with very short $\Delta t$. Long $\Delta t$’s coming from unaccounted for dead time are seen on the right of this plot. Error bars represent statistical uncertainties.
Figure 7: $\chi^{2}$/NDF distribution of Erlang fit for multiple-muon data in weekly binning. Since the dominant source of selected events are cosmic ray multiple muon showers following Poissonian arrival times, the Erlang fit used to extract the rate is well matched to the data and robust over time.
The principal advantage of this method is its robustness to irregularities in the data stream. Occasional detector dead times or DAQ interruptions can produce unusually large $\Delta t$ values. In the $\log_{10}(\Delta t)$ histogram, such events populate the far tail of the distribution and lie well outside the region around the peak that is used in the fit. Similarly, very small $\Delta t$ values can arise from split tracks or double triggers in the DAQ. These appear well below the main peak and are again excluded from the fitting region. As a result, the fitted event rate is determined only from the central part of the distribution, where the data follow the expected exponential law that comes from the Poissonian process of the signal (multiple muons). This allows calculation of error bars on the rates from pure statistics, by eliminating systematic errors in both the numerator and denominator of the naïve “divide count by the live time of a file” process used in Acero et al. (2019).
@@
Figure 8 shows the variation of the multiple-muon event rate seen in those four years of ND data, as well as the effective temperature. Both the data and the temperature are presented in weekly bins. The percentage variation is calculated using
@@
The muon rate in the multiple-muon data displays a seasonal trend that is out of phase with the effective atmospheric temperature. Although periodic, there is no a priori reason to expect a cosine shape. However, it does serve to extract the phase of the magnitude maxima, so the percentage variations of multiple-muon rate and effective temperature are fitted with cosine curves,
@@
where $V_{0}$ is the average rate, $V$ is the magnitude, $t_{year}$ is the time period (fixed at one year) and $\phi$ is the phase. The fitted parameters are shown in Table 1. The multiple-muon rate varies by $\pm 2\%$ in rate, and has the opposite phase as the temperature. The correlation coefficient between the rate of multiple muons and the effective temperature is found to be $0.98\pm 0.03$.
Table 1: Best fit parameters for the cosine fit of percentage variation of the multiple-muon data and effective temperature
@@
VI Explanation
VI.1 The altitude-geometry explanation
The total mass of the atmosphere remains approximately constant throughout the year, but seasonal temperature differences cause vertical expansion of atmosphere in summer and contraction in winter. Under the assumption of an isothermal atmosphere, the ideal gas law $PV=nRT$ implies that a $\pm$2% change in absolute temperature results in a corresponding $\pm$2% shift in altitude for a given pressure level. As shown in Fig. 9 Kokoulin et al. (2016), for a shower origin at fixed pressure, the increased altitude (point F) in summer leads to a broader lateral spread of muons at the detector. In contrast, during winter, the lower altitude (point E) results in a more concentrated footprint.
Figure 10: Production altitude distribution of multiple-muon events from CORSIKA with muon threshold energy of $54\text{\,}\mathrm{GeV}$. Three winter months (December–February) and three summer months (June–August) are compared. Vertical lines represent the mean muon birth altitude for summer (red dashed line) and winter (black solid line).
Figure 11: Muon separation distribution of multiple-muon events from CORSIKA with muon threshold energy of $54\text{\,}\mathrm{GeV}$. Three winter months (December–February) and three summer months (June–August) are compared. Vertical lines represent the mean muon separation for summer (red dashed line) and winter (black solid line).
We use the events generated with the CORSIKA simulation to study the altitude and muon separation variation during summer and winter. Figure 10 shows the mean birth altitude of muons in the upper atmospehere from multiple muon events is higher in summer than in winter. As a result, the separation among muons for those events also increases in summer compared to winter, as shown in the Fig. 11. It is instructive to consider the typical transverse momentum of mesons in hadronic showers, and how that impacts the separation distribution of muons reaching a detector. The opening angle between a secondary pion and the primary cosmic ray direction in an air shower is approximately proportional to $p_{t}/p_{l}$, where $p_{t}$ is the pion’s transverse momentum and $p_{l}\sim E_{\pi}\sim E^{\mu}$ is its longitudinal momentum. Consequently, the typical lateral separation between the two muons originating from different pions in the same shower can be expressed as
$\sqrt{2}\times h\times(p_{t}/E^{\mu}),$ (10)
where $h$ is the altitude at which the pions decay. Since only higher-energy muons can reach deeper underground detectors, the required $E^{\mu}$ increases with depth, leading to smaller lateral separations between muons for such underground detectors. To illustrate how the altitude-geometry effect depends on detector depth and muon energy, Table 2 shows the typical lateral separation of muons at various underground depths, calculated using Eq. 10 for pions decaying at an altitude of 20 km and a typical pion $p_{t}$ of $300~\text{MeV/c}$. The corresponding minimum muon energies to reach the detector are calculated assuming standard rock with a density of $2.8~\text{g/cm}^{3}$ and a minimum ionizing energy loss of $2~\text{MeV}/(\text{g}\,\text{cm}^{-2})$. The muon separation decreases with increasing the energy (required to reach deeper detectors) and with increasing depth due to energy-dependent selection of muons closer to the shower core.
Table 2: Typical lateral separation of muons as a function of underground detector depth and muon energy threshold, calculated using Eq. 10. The separation indicates the typical transverse spread of muons from a cosmic ray shower at the detector level.
Depth $E^{\mu}_{min}$ Muon
(m) (GeV) Separation (m)
18 10 850
80 50 170
180 100 85
800 500 17
1800 1000 8.5
For the NOvA or MINOS underground near detectors, with an energy threshold close to $50\text{\,}\mathrm{G}\mathrm{e}\mathrm{V}$, the typical muon separation is $170\text{\,}\mathrm{m}$ which is much larger than the detector size. We can use Table 2 to determine whether any underground detector with a particular depth and area ($x\times z$) exhibits a winter or summer maximum. If the separation is much smaller than $x$ or $z$, a summer maximum will be seen. If the separation is larger or comparable, a winter maximum will be observed. However, the single muon rate remains unaffected by the altitude-geometry effect. For every single muon that misses a finite-sized detector because it starts at a higher altitude in the summer, another muon that would miss the detector in the summer now hits it. The muon flux does depend on the detector geometry.
The role of summer/winter differences in the origin altitude of muons in an air shower, and its effect on shower containment in a finite detector, was also considered before by MINOS Adamson et al. (2015). However, the conclusion in the paper was based on the distributions of observed muon separation in the detectors, instead of the true separation between muons. That reference based its conclusion on the measured muon separation distribution, assuming that it was the true distribution.
The dependence of the altitude–geometry effect as function of detector depth can also explain the seasonal behaviors observed in other experiments. For deeper detectors that require higher muon energy (TeV) thresholds, the altitude–geometry effect is minimal. The surface area of the MACRO experiment Ambrosio et al. (2002) is $76\text{\,}\mathrm{m}$$\times$$12\text{\,}\mathrm{m}$ and the threshold energy required for an atmospheric muon to reach the detector is $1.3\text{\,}\mathrm{T}\mathrm{e}\mathrm{V}$. The average separation among muons in this case, is smaller than the size of the detector. Hence, MACRO observed a summer maximum for multiple-muon events Ronga (2017). In the MINOS FD Adamson et al. (2015), the typical muon separation from an air shower is comparable to the size of the detector. So, the observed muon separation determines whether there is a summer maximum or a winter maximum. Calculations by members of the IceCube collaboration based on the altitude-geometry effect seemed to explain the MINOS and NOvA multiple-muon seasonal data Gaisser and Verpoest (2021a, b). The altitude-geometry effect was considered responsible for the seasonal results of the DECOR experiment Yurina et al. (2019), which measured muons close to the surface of the Earth. For those muons, with an energy typically of a few GeV, other effects such as pressure changes and muon decay are important, as discussed further in Sec. VI.2.
VI.2 Other seasonal explanation
The MINOS result in Ref. Adamson et al. (2015) disfavored dimuon decays or temperature effects as contributing to the explanation of multiple-muon variation. However, it considered the anticorrelation between the primary meson interaction and decay as the most plausible explanation for the observed phenomenon. According to this hypothesis, higher temperatures in summer increase the decay probability of mesons, leaving fewer mesons available to inelastically scatter and produce additional secondaries in subsequent generations of the hadronic shower. To test this independently of CORSIKA, we used a simulation of hadronic showers in the atmosphere with densities corresponding to summer and winter Guan and Goodman (2023). We followed the meson and muon content in the hadronic shower as subsequent generations of both mesons and nucleons interacted, as a function of primary cosmic ray energy. The simulation showed that a shower produces more muons when the primary meson decays rather than interacts, contrary to the hypothesis favored by MINOS Adamson et al. (2015). In addition, this explanation would predict a winter maximum in an infinite detector, in contradiction with the CORSIKA result in Fig. 1.
@@ VI.2 Other seasonal explanation
Table 3: Analytic calculation of the percentage of muons that reach sea level before they decay as a function of muon energy. $E_{\mu}$ is the muon energy, $N_{0}$ is the initial number of muons, and $N$ is the number of muons that survive. The summer/winter altitude difference is assumed to be 2%.
@@
Another potential explanation for the observed seasonal pattern of multiple muons is muon decay, which was not discussed in Ref. Adamson et al. (2015). Due to relativistic effects, muon decay is more significant at lower muon energies, and thus its impact is expected to be largest for detectors located at or near the surface. An analytical estimate of the seasonal modulation due to muon decay for various energy thresholds is summarized in Table 3. Surface detectors such as the NOvA FD Acero et al. (2021), DECOR Yurina et al. (2019); Tolkacheva et al. (2011), and GRAPES Arunbabu et al. (2017), all with muon thresholds near 5 GeV, can exhibit a 0.6% excess of events during the winter. In contrast, detectors such as the NOvA ND and MINOS ND, which have muon thresholds close to 50 GeV, would only show a smaller 0.09% excess in winter, which is much smaller than the observed winter excess, and also inconsistent with Fig. 1. Muon decay is even less important for the MINOS FD and MACRO.
VII Conclusion
@@
The altitude-geometry effect is the most relevant explanation for the opposite seasonal behaviors between single and multiple-muon events. Evidence supporting this conclusion comes from a comparison between the typical muon separation distribution as a function of underground detector depth vis-à-vis the detector size, using the CORSIKA simulation, which demonstrates a summer maximum for multiple-muon events in an infinite detector. However, for the finite-sized NOvA or MINOS ND, the simulation shows a winter maximum for multiple-muon events. The effect of muon decay for $54\text{\,}\mathrm{GeV}$ muons was found to be too small. In general, the altitude-geometry effect leads to a winter maximum for multiple muons in underground detectors if the muon separation is similar to or larger than the size of the detector, but when it is much smaller than the detector size, a summer maximum is to be expected. This agrees with observations from the MINOS FD, MINOS ND, NOvA ND, and MACRO. In particular, the results of Ref. Adamson et al. (2015) and Acero et al. (2019) are now understood.